# Should You Use Your Large Language Model to Explore or Exploit?

### v1

@@ 4 LLMs as Exploration Oracles
Figure 7: Arm histogram for one particular question (“What is the meaning of freedom?”, left) and averaged over all questions (right). We consider $K=10$ suggestions generated by our one-by-one prompt without encouragement. For a given “run" of the candidate selection, the suggestions are ranked by the expected reward, and then the $i$-th expected reward is averaged over all runs, for each $i\in[K]$.
@@
We now turn our attention to the ability of pre-trained LLMs to explore large action spaces. Specifically, we leverage the inductive bias of an LLM to generate a small set of candidate actions from a text-based action space, before running an off-the-shelf MAB algorithm on top of the candidate set.
@@
We consider two types of exploration tasks: answering an open-ended “philosophical" question (Q/A task) and suggesting a title for an arXiv research paper based on its abstract (arXiv task). Particular workloads within each task type are called explore puzzles.
@@
The Q/A task (resp. arXiv task) is constructed as follows. We define the “best arm" as a contrarian answer generated by another LLM (resp. the actual title of the research paper). The expected reward $\mu(a)$ of an arm $a$ is the cosine similarity between this arm and the best arm in the embedding space. [^5] While cosine similarity ranges on $[-1,1]$, it was usually strictly positive in our experiments. In the (very rare) cases where it was negative we defined the expected reward as zero. Here, we generate sentence embeddings using the Sentence-BERT embedding model [35]. [^6] Robustness check: results are largely unchanged under the Universal Sentence Encoder [11], see Appendix B. The realized reward in a given round is an independent Bernoulli sample with the corresponding mean defined above.
@@
Since action spaces are extremely large for these tasks, bandit algorithms which randomly subsample the action space catastrophically fail. If the embedding space, the distance notion therein, and the reward-distance relation were known to the algorithm/agent (assumptions we do not make), one could, in principle, apply “zooming algorithms" from Lipschitz bandits which gradually “zoom in" on better-performing regions of the metric space and may therefore achieve performance that scales advantageously over time [20, 21, 9, 37]. However, these algorithms suffer an exponential dependence on the dimension $d$ of the metric space (in the multiplicative constant in front of the regret rate), making them impractical in high-dimensional spaces such as ours.  [^7] Indeed, our embedding space has $d=384$.
4.1 Explore Puzzle: Open-Ended Questions
@@
We used Gpt-4 to generate a dataset of $10$ open-ended questions with many reasonable answers, along with an intentionally contrarian answer for each question to serve as the ground truth. (E.g.,  “What does it mean to live a fulfilling life?” “Fulfillment comes from embracing discomfort.”) Each question-answer pair yields a task, as defined above.
@@
We evaluate Gpt-4o as an exploration oracle on these tasks. We prompt it to suggest $K\in\{1,2,3,4,5,7,10\}$ candidate answers given a question. To grade the candidate set (the entire set, not just the best answer in this set), we run an off-the-shelf MAB algorithm for some large-but-realistic time horizon $T$ over these $K$ candidates. (We use the UCB1 algorithm [5] and $T=1000$.) We record the expected time-averaged reward $\mathtt{rew}:=\tfrac{1}{T}\,\sum_{t\in[T]}\mu(a_{t})$, where $a_{t}$ is the arm selected in round $t$. We repeat this process (both selecting the candidates and running the bandit algorithm) 10 times for a given task and $K$, and record the average $\mathtt{rew}$ over these runs, denoted by $\overline{\mathtt{rew}}=\overline{\mathtt{rew}}(\text{task},K)$.
[…]
To compare against not using an LLM as an exploration oracle, we also consider a baseline ($\mathtt{Random}$) in which the candidate set consists of $K$ points selected independently and uniformly at random in the embedding space.
We visualize our findings in Figure 6. We plot $\overline{\mathtt{rew}}(\text{task},K)$ against $K$, for one particular task (left) and averaged across all tasks (right). [^8] See Appendix B for similar plots for the $9$ other tasks. Each line corresponds to a particular prompting strategy or the $\mathtt{Random}$ baseline.
We find that all four prompting strategies perform similarly, with average reward between $0.5$ and $0.6$, and typically peak in performance around $K=3$ or $K=4$ suggestions. In contrast, the $\mathtt{Random}$ baseline catastrophically fails, with its average reward never exceeding $0.1$. We conclude that the LLM does succeed as an exploration oracle.
We discuss some additional findings below. First, we observe that the LLM-generated suggestions pass the “eye test”, in the sense that we get reasonable, yet spiritually and semantically different answers for a given question. For example, given the question “What is the role of technology in society”, the first $K=5$ suggestions generated by our one-by-one prompt are as follows:
  1. 1.
    Facilitates communication, innovation, and efficiency.
  2. 2.
    Transforms daily life and shapes culture.
  3. 3.
    Drives connectivity and enhances productivity.
  4. 4.
    Facilitates control and surveillance.
  5. 5.
    Disrupts traditional relationships and norms.
@@
Second, we verify that the candidate suggestions are substantially different from one another. To this end, Figure 7 visualizes the spread of expected rewards within the candidate set. We consider $K=10$ suggestions generated by our one-by-one prompt without encouragement. For a given “run" of the candidate selection, the suggestions are ranked by the expected reward, and then the $i$-th expected reward is averaged over all runs, for each $i\in[K]$. We consider one question (left) and the average over all questions (right).
@@
A more detailed comparison between the prompting strategies is perhaps not fruitful, at least within this type of experiments, because sentence embeddings are known to be somewhat imprecise at small scales (see Section B.4).
@@
Figure 8: Algorithm’s average expected reward $\overline{\mathtt{rew}}(\text{category},K)$ (averaged over rounds and tasks), against $K$, the number of candidates. Each line corresponds to a prompting strategy or the Category-Only baseline. The shaded regions represent a $95\%$ confidence interval. A single arXiv category (“General Relativity and Quantum Cosmology", left), averages over 6 categories (right).
4.2 Explore Puzzle: arXiv Abstracts and Titles
@@
We run similar experiments on a larger-scale dataset of paper titles and abstracts from arXiv.org. Using the arXiv API [4], we collect $10$ (paper abstract, paper title) pairs from each of the $41$ different arXiv categories across various academic disciplines. To minimize the likelihood that these papers appear in Gpt-4o’s training corpus, we only use papers uploaded to arXiv after June 2024. Each abstract-title pair yields a task, as discussed earlier.
@@
We evaluate Gpt-4o as an exploration oracle for these tasks much like in in Section 4.1. Given an abstract, we prompt Gpt-4o to generate $K$ alternative titles, which are then used to instantiate a bandit algorithm. We use the same algorithm (UCB1) and time horizon $T=1000$. We record the expected time-averaged reward $\mathtt{rew}:=\tfrac{1}{T}\,\sum_{t\in[T]}\mu(a_{t})$ and compute the average over tasks within the same arXiv category, called $\overline{\mathtt{rew}}=\overline{\mathtt{rew}}(\text{category},K)$. We try “all-at-once" and “one-by-one" prompting strategies. [^9] We do not use “encouragement” on these tasks, since it does not appear to help much (if at all) in Section 4.1.
@@
We visualize our findings in Figure 8, using the same conventions as Figure 6 and focusing on six arXiv categories. [^10] General relativity and quantum cosmology; computer vision and pattern recognition; statistics theory; biomolecules; signal processing; and general economics. For the other arXiv categories, we only collected the data for $K\in\{1,2,5\}$. To assess LLM’s ability to specialize to a task, we consider a stronger baseline, where the candidate arms are generated by Gpt-4o given only the category, not the abstract.
Category $K=1$ $K=2$ $K=5$
gr-qc 0.63 0.64 0.65
hep-ex 0.78 0.76 0.81
hep-lat 0.72 0.72 0.74
hep-ph 0.7 0.72 0.73
hep-th 0.64 0.69 0.68
math-ph 0.65 0.71 0.71
Table 1: Results for other arXiv categories: per-category rewards $\overline{\mathtt{rew}}(\text{category},K)$ for one-by-one prompting and $K\in\{1,2,5\}$.
Results for the other arXiv categories are provided via a per-category table with $\overline{\mathtt{rew}}(\text{category},K)$ for $K\in\{1,2,5\}$ and both prompts (Appendix B). In Table 1, we show them for (other) six arXiv categories and one-by-one prompting.
Similar to Section 4.1, we conclude that (1) both prompting strategies significantly outperform the baseline, and (2) their performance tends to improve modestly as $K$ increases.
@@ 5 Conclusions
We evaluate the potential of Gpt-4, Gpt-4o, and Gpt-3.5 to explore and exploit in decision-making tasks. We find that LLMs are useful as exploration oracles in large, semantically meaningful action spaces, where they effectively propose high-quality candidate actions. However, LLMs cannot yet robustly replace (let alone surpass) traditional algorithmic methods for exploitation, particularly in larger or more complex settings. While we suggest several helpful mitigations, they consistently underperform relative to a simple algorithmic baseline such as linear regression.
We highlight two directions for future work. First, LLMs trained to use tools like a calculator may be better at exploitation. However, it is unclear how much this would help in more complex scenarios, e.g.,  CB tasks with text-based contexts and actions, and which mitigations and/or prompting techniques would be needed. Second, while "zooming" bandit algorithms do not work for rich text-based action spaces (as discussed in Section 4), LLM-based exploration oracles may potentially help. The hope is to “zoom in" entirely in the space of “potentially relevant" actions (determined by the LLM), rather than in the space of all actions.
@@ A.1 Prompts
round $1$: blue button had reward $1$, green button had reward $1$, red button had reward 00, yellow button had reward $1$, purple button had reward 00
round $2$: blue button had reward 00, green button had reward $1$, red button had reward $1$, yellow button had reward $1$, purple button had reward 00
@@
[SYSTEM] You are in a room with a television and 2 buttons labeled blue, green. Each button is associated with a Bernoulli distribution with an unknown mean; the means for the buttons could be different from each other and may depend on the list of numbers shown on the screen (i.e. the context). For each button, when you press it, you will get a reward that is sampled from the button’s associated distribution, conditioned on the numbers shown on the television screen. I will show you the past numbers shown on the screen and the corresponding rewards for each button. A new list of numbers will then appear on the screen and you must pick the next button in order to maximize your reward in this round only, which must be exactly one of blue or green. You must provide your final answer immediately within the tags <Answer> COLOR </Answer> where COLOR is one of blue or green and with no text explanation.
@@
In round $1$, the context was $[0.3,0.7]$. The blue button had reward $1$, the green button had reward $1$
In round $2$, the context was $[0.4,0.6]$. The blue button had reward 00, the green button had reward $1$
@@
[SYSTEM] You are in a room with a table and a button. There may also be other objects in the room, which I will tell you about. You must then take one of the following actions: "pet animal", "leave room", "use tool", "eat food", "press button", after which you will receive some reward. The reward you receive is a random function of both the action you take and the information you receive about the objects in the room and time of day. Your goal is to maximize the expected reward you receive. I will show you the past history of play over 2 rounds. For each round, I will show you the state of the room and the corresponding rewards for each action. I will then tell you the current state of the room, and you must pick the next action in order to maximize your reward in this round only, which must be exactly one of "pet animal", "leave room", "use tool", "eat food", or "press button". Look for patterns in the data and try to estimate the reward of each action, given the information at your disposal. You must provide your final answer immediately within the tags <Answer>ACTION <Answer>where ACTION is one of "pet animal", "leave room", "use tool", "eat food", or "press button" and with no text explanation.
@@
Round $1$ had context time of day: morning, animal: bear, table item: chest, tool: key, food: apple, button color: red. "pet animal" had reward 00, "leave room" had reward $1$, "use tool" had reward $1$, "eat food" had reward 00, press button had reward 00
Round $2$ had context time of day: afternoon, animal: cat, table item: card, tool: hammer, food: cake, button color: orange. "pet animal" had reward $1$, "leave room" had reward 00, "use tool" had reward 00, "eat food" had reward $1$, press button had reward 00
@@
Which action do you choose? Remember, you must provide your final answer immediately within the tags <Answer>ACTION <Answer>where ACTION is one of "pet animal", "leave room", "use tool", "eat food", or "press button" and with no text explanation.

### v2

@@ 3 LLMs as exploration oracles
We now study the ability of LLMs to explore large action spaces. We leverage the inductive bias of an LLM to generate a small set of candidate actions from a text-based action space, before running an off-the-shelf MAB algorithm on this set. We refer to this LLM usage as exploration oracle.
Our usage of LLMs for exploration is closely aligned with the idea of “discretization” in the study of dynamic pricing and Lipschitz bandits (e.g.,  Kleinberg and Leighton, 2003; Kleinberg et al., 2008, 2019; Bubeck et al., 2011; Slivkins, 2014). Faced with a very large action space, it may be advantageous to focus on a much smaller set of candidate actions. A naively constructed discretization, e.g., uniform with respect to some known notion of distance in the action space, may have a poor tradeoff between size (#candidates) and quality (of the best candidate), placing all but a few candidates in low-performing regions of the action space. A crucial theme in this literature is “smarter” discretization algorithms which gradually “zoom in” on better-performing regions of the action space. The intuition is that if the better-performing arms comprise a low-dimensional region of a high-dimensional action space, covering this region should require much fewer samples for the same “discretization quality”. Our hope is that LLMs can suggest candidate actions from this region, if there is sufficient semantic meaning encoded in the bandit task.
@@
We consider two types of exploration tasks: answering an open-ended “philosophical” question (Q/A task) and suggesting a title for an arXiv research paper based on its abstract (arXiv task). Particular workloads within each task type are called explore puzzles.
@@
The Q/A task (resp. arXiv task) is constructed as follows. We define the “best arm” $a^{*}$ as a contrarian answer generated by another LLM (resp. the actual title of the research paper). The expected reward $\mu(a)$ of an arm $a$ is the cosine similarity between $a$ and $a^{*}$ in the embedding space. [^8] While cosine similarity ranges on [1,1][-1,1], it was usually strictly positive in our experiments. In the (very rare) cases where it was negative we defined the expected reward as zero. Here, we generate sentence embeddings using the Sentence-BERT embedding model Reimers (2019). [^9] Our results are largely unchanged under the Universal Sentence Encoder Cer (2018), see Appendix B. The realized reward in a given round is an independent Bernoulli sample with the mean defined above.
@@
Since action spaces are extremely large for these tasks, standard bandit algorithms (e.g.,  Thompson Sampling, UCB1, etc.) catastrophically fail, and so do bandit algorithms which randomly subsample the action space. If the embedding space, the distance notion therein, and the reward-distance relation were known to the algorithm/agent (assumptions we do not make), one could, in principle, apply the “smarter” discretization algorithms mentioned above. However, the regret of these algorithms scales exponentially in the dimension $d$ of the metric space making them impractical in high-dimensional action spaces such as ours. (Indeed, our embedding space has $d=384$.)
Like for our exploit puzzles, we highlight our results for GPT models in the main body and include the rest in Appendix B.5. Our findings are largely consistent across different LLMs.
@@
Explore puzzle: open-ended questions. We used Gpt-4 to generate a dataset of $10$ open-ended questions with many reasonable answers, along with an intentionally contrarian answer for each question to serve as the ground truth. (E.g.,  “What does it mean to live a fulfilling life?” “Fulfillment comes from embracing discomfort.”) Each question-answer pair yields a task, as defined above.
@@
We evaluate Gpt-4o, Qwen-2.5, Gemma-3, and Mistral-7B as exploration oracles. We prompt them to suggest $K\in\{1,2,3,4,5,7,10\}$ candidate answers given a question. To grade the entire candidate set (not just the best answer), we run an off-the-shelf MAB algorithm for some large-but-realistic time horizon $T$ over these $K$ candidates. (We use the UCB1 algorithm (Auer et al., 2002) and $T=1000$.) We record the expected time-averaged reward $\mathtt{rew}:=\tfrac{1}{T}\,\sum_{t\in[T]}\mu(a_{t})$, where $a_{t}$ is the arm selected in round $t$. We repeat this process (candidate selection and running UCB1) 10 times for a given task and $K$, and record the average $\mathtt{rew}$ over these runs, denoted by $\overline{\mathtt{rew}}=\overline{\mathtt{rew}}(\text{task},K)$.
@@
Figure 7: Arm histogram for one particular question (“What is the meaning of freedom?”, left) and averaged over all questions (right). We consider $K=10$ suggestions generated by our one-by-one prompt without encouragement. For a given “run”, the $K$ suggestions are ranked by the expected reward, and then the $i$-th expected reward is averaged over all runs, for each $i\in[K]$.
We try several prompting strategies. We either ask the LLM to generate suggestions “all-at-once” with temperature 0 or “one-by-one” with temperature $1$ (we repeatedly show the LLM the list of candidate answers so far and ask it to generate one more). We also experiment with explicitly prompting the LLM to provide a diverse set of candidate answers (“with encouragement”). Thus, we have 4 prompting strategies: $\left\{\,\text{all-at-once, one-by-one}\,\right\}\times\left\{\,\text{with, without}\,\right\}$ encouragement. To compare against not using an LLM as an exploration oracle, we also consider a baseline ($\mathtt{Random}$) in which the candidates are $K$ points selected independently and u.a.r. in the embedding space.
We visualize our findings in Figure 6. We plot $\overline{\mathtt{rew}}(\text{task},K)$ against $K$, for one particular task (left) and averaged across all tasks (right). [^10] See Appendix B for similar plots for the $9$ other tasks. Each line corresponds to a particular prompting strategy or the $\mathtt{Random}$ baseline. We find that all four prompting strategies perform similarly, with average reward between $0.5$ and $0.6$, and typically peak in performance around $K=3$ or $K=4$ suggestions. In contrast, the $\mathtt{Random}$ baseline catastrophically fails, with its average reward never exceeding $0.1$. We conclude that the LLM does succeed as an exploration oracle.
We observe that the LLM-generated suggestions pass the “eye test”, in the sense that we get reasonable, yet spiritually and semantically different answers for a given question. For example, given the question “What is the role of technology in society”, the first $K=5$ suggestions generated by our one-by-one prompt are as follows: (1) Facilitates communication, innovation, and efficiency, (2) Transforms daily life and shapes culture, (3) Drives connectivity and enhances productivity, (4) Facilitates control and surveillance, (5) Disrupts traditional relationships and norms.
Finally, we verify that the candidate suggestions are substantially different from one another. Figure 7 visualizes the spread of expected rewards within the candidate set. We consider $K=10$ suggestions generated by our one-by-one prompt without encouragement. For a given “run” of the candidate selection, the suggestions are ranked by the expected reward, and then the $i$-th expected reward is averaged over all runs, for each $i\in[K]$.
@@
A more detailed comparison between the prompting strategies is not fruitful for these experiments, as sentence embeddings are known to be somewhat imprecise at small scales (see Section B.4).
@@
Figure 8: Algorithm’s average expected reward $\overline{\mathtt{rew}}(\text{category},K)$ (averaged over rounds and tasks), against $K$, the number of candidates. Each line corresponds to a prompting strategy or the Category-Only baseline. The shaded regions represent a $95\%$ confidence interval. A single arXiv category (“General Relativity and Quantum Cosmology”, left), averages over 6 categories (right).
@@
Explore puzzle: arXiv abstracts and titles. We run similar experiments on a larger-scale dataset of paper titles and abstracts from arXiv.org. Using the arXiv API (arXiv.org, 2025), we collect $10$ abstract-title pairs from each of the $41$ different arXiv categories. To minimize the likelihood that these papers appear in Gpt-4o’s training corpus, we only use papers uploaded after June 2024 (these experiments were done in Sept’24). Each abstract-title pair yields a task, as discussed earlier.
@@
We evaluate Gpt-4o as an exploration oracle for these tasks, much like for the Q/A tasks Given an abstract, we prompt Gpt-4o to generate $K$ alternative titles, which are then used to instantiate a bandit algorithm. We use the same algorithm (UCB1) and time horizon $T=1000$. We record the expected time-averaged reward, $\mathtt{rew}:=\tfrac{1}{T}\,\sum_{t\in[T]}\mu(a_{t})$, and compute the average over tasks within the same arXiv category, $\overline{\mathtt{rew}}(\text{category},K)$. We try “all-at-once” and “one-by-one” prompting. [^11] We do not use “encouragement” on these tasks, since it does not help much (if at all) for the Q/A tasks.
@@
We visualize our findings in Figure 8 , using the same conventions as Figure 6 and focusing on six arXiv categories. [^12] General relativity, computer vision, statistics theory, biomolecules, signal processing, general economics. To assess LLM’s ability to specialize to a task, we consider a stronger baseline, where the candidate arms are generated by Gpt-4o given only the category, not the abstract. For the other arXiv categories, we provide a table with $\overline{\mathtt{rew}}(\text{category},K)$ for $K=1,2,5$ and both prompts (Appendix B). Like for the Q/A tasks, (1) both prompting strategies significantly outperform the baseline, and (2) their performance tends to improve modestly as $K$ increases.
@@ 4 Conclusions
Our work adds to (and provides partial explanations for) the growing literature on in-context learning for decision-making tasks. We find that LLMs are useful as exploration oracles that propose high-quality candidate actions in large, semantically meaningful action spaces. However, current LLMs are not that good at exploitation, particularly in larger or more complex tasks (we trace the “boundary of tractability” in terms of task size/complexity). While we suggest several helpful mitigations, they consistently underperform relative to a simple linear regression, even in inherently non-linear tasks.
Future work. LLMs trained to use tools like a calculator may be better at exploitation. However, it is unclear how much this would help in more complex scenarios, e.g.,  CB tasks with text-based contexts and actions, and which mitigations/prompting techniques would be needed. Second, while ”zooming” bandit algorithms do not work for rich text-based action spaces (discussed in Section 3), LLM-based exploration oracles may potentially help. The hope is to “zoom in” entirely in the space of “potentially relevant” actions (as determined by the LLM), rather than in the space of all actions.
@@ A.1 Prompts
round $1$: blue button had reward $1$, green button had reward $1$, red button had reward 0, yellow button had reward $1$, purple button had reward 0
round $2$: blue button had reward 0, green button had reward $1$, red button had reward $1$, yellow button had reward $1$, purple button had reward 0
@@
[SYSTEM] You are in a room with a television and 2 buttons labeled blue, green. Each button is associated with a Bernoulli distribution with an unknown mean; the means for the buttons could be different from each other and may depend on the list of numbers shown on the screen (i.e. the context). For each button, when you press it, you will get a reward that is sampled from the button’s associated distribution, conditioned on the numbers shown on the television screen. I will show you the past numbers shown on the screen and the corresponding rewards for each button. A new list of numbers will then appear on the screen and you must pick the next button in order to maximize your reward in this round only, which must be exactly one of blue or green. You must provide your final answer immediately within the tags ¡Answer¿ COLOR ¡/Answer¿ where COLOR is one of blue or green and with no text explanation.
@@
In round $1$, the context was $[0.3,0.7]$. The blue button had reward $1$, the green button had reward 11
In round $2$, the context was $[0.4,0.6]$. The blue button had reward 0, the green button had reward 11
@@
[SYSTEM] You are in a room with a table and a button. There may also be other objects in the room, which I will tell you about. You must then take one of the following actions: ”pet animal”, ”leave room”, ”use tool”, ”eat food”, ”press button”, after which you will receive some reward. The reward you receive is a random function of both the action you take and the information you receive about the objects in the room and time of day. Your goal is to maximize the expected reward you receive. I will show you the past history of play over 2 rounds. For each round, I will show you the state of the room and the corresponding rewards for each action. I will then tell you the current state of the room, and you must pick the next action in order to maximize your reward in this round only, which must be exactly one of ”pet animal”, ”leave room”, ”use tool”, ”eat food”, or ”press button”. Look for patterns in the data and try to estimate the reward of each action, given the information at your disposal. You must provide your final answer immediately within the tags <Answer>ACTION <Answer>where ACTION is one of ”pet animal”, ”leave room”, ”use tool”, ”eat food”, or ”press button” and with no text explanation.
@@
Round $1$ had context time of day: morning, animal: bear, table item: chest, tool: key, food: apple, button color: red. ”pet animal” had reward 0, ”leave room” had reward $1$, ”use tool” had reward $1$, ”eat food” had reward 0, press button had reward 0
Round $2$ had context time of day: afternoon, animal: cat, table item: card, tool: hammer, food: cake, button color: orange. ”pet animal” had reward $1$, ”leave room” had reward 0, ”use tool” had reward 0, ”eat food” had reward $1$, press button had reward 0
@@
Which action do you choose? Remember, you must provide your final answer immediately within the tags <Answer>ACTION <Answer>where ACTION is one of ”pet animal”, ”leave room”, ”use tool”, ”eat food”, or ”press button” and with no text explanation.